1. Janet wants to put a wallpaper border around her rectangular kitchen. The room is 12 by 20 feet a. What is the total length of wallpaper border she needs? b. If the wallpaper border comes in packages of 12 feet, how many packages must she buy? c If the packages are priced at $11.87 each, what will be the total cost of the wallpaper? a. What is the total length of wallpaper border needed? feet (Simplify your answer.) b. How many 12 ft packages are needed to finish the project? packages (Simplify your answer.) c. What is the total cost of the border? 2. Translate the sentence to an equation and solve it. One -third of a number is seven-twelfths What is the equation that represents the sentence? The solution is X 3. Translate to an inequality and then solve. Sixteen subtracted from a number is greater than thirty-nine. Translate into an inequality, ?
Answer #1

Concept: Perimeter of rectangle.

Then finding number of packages and converting it into next whole number.

Then finding total cost.

D- 12 64 12. = 5-33 87 c) ll cost ef one package $ 11.87 ome ackag Total Cost= 11-87X6 11 22